prove that 1+2 sin2=0cos20
                  1+3 tan20
prove this sir i am not getting this

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Tauseef
 
LHS
1 + sin² 0° 
Now as  sin 0° = 0
 
Hence
1 + 0 = 1 = cos ² 0°
 
Now
 
1 + 3 tan²0°
 = 1 + 0 = 1 = cos² 0°
 
 
 
Regards
Arun (askIITians forum expert)